Gary is x years old. Sarah is 4 years older. Kevin is three times Sarah's age. The total of all their ages is 76 years. Create an equation and then solve it to find out how old each person is.

Answer by TimothyLamb(4379)

x = gary
y = sarah
z = kevin
---
y = x + 4
z = 3y
x + y + z = 76
---
put the system of linear equations into standard form
---
x - y = -4
3y - z = 0
x + y + z = 76
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = gary = 12
y = sarah = 16
z = kevin = 48
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php